# Advancing Generative AI for Portuguese with Open Decoder Gervásio PT*

### 29 Feb 2024 v1

@@ 1.   Introduction
By exploiting this approach of continuing the training of a previous strong foundation model, we contribute a new model with instruction tuning to foster the technological preparation of the Portuguese language. To the best of our knowledge, this is the first decoder under the Transformer architecture that is both (i) specifically improved for Portuguese, covering two variants of this language, namely PTBR and PTPT, and (ii) fully open, that is it cumulatively complies with all the features of being open source and openly distributed for free under a most permissive license (including for research and for commercial purposes). The model is available at [https://huggingface.co/PORTULAN].
To the best of our knowledge and at the time of writing, Gervásio represents the state of the art reported in the literature for open, 7 billion parameter decoders for Portuguese, surpassing the model it is based on as well as otherdecoders for Portuguese of similar size. The release of Gervásio, alongside the instruction datasets used to train it and which are also a novel contribution of this paper, seeks to contribute to foster research and innovation for the language technology for Portuguese.
@@ 2.1.   Decoders for Portuguese
Looking for decoders specifically developed or improved for Portuguese that are publicly distributed and for which it is possible to find a publicly available report, to the best of our knowledge there can be found only two that, with 7 billion parameters or more, match or surpass the size of Gervásio PT* contributed in the present paper, namely the Sabiá models with 7 and 65 billion parameters (Pires et al., 2023) It is worth noting that: these two models were developed for only one of the variants of Portuguese, PTBR, but not for PTPT; the 65 billion parameter model is reported in that publication but it is not distributed; and the 7 billion parameter model is distributed in a non open license, being its reuse restricted to research purposes only.
Other decoders that at the time of writing the present paper, can be found of comparable size are not documented, besides being also only for only one of the variants of Portuguese, namely PTBR: Boana, Cabra, Cabrita, Canarim. [^1] All on HuggingFace, at lrds-code/boana-7b-instruct, nicolasdec/CabraMistral7b-0.2, 22h/open-cabrita3b, and dominguesm/canarim-7b, respectively.
The other decoders, numbering about a dozen, that can be found for Portuguese have a smaller size, and are also only for PTBR. The largest of these, with 3 billion parameters and named Cabrita 3B, is distributed in Hugging Face (HF) and documented in a non peer-reviewed publication (Larcher et al., 2023). The second largest is Aira [^2] https://huggingface.co/nicholasKluge/Aira-2-portuguese-1B7 and it is presented in HF as having 1.7 billion parameters and being based on Bloom. No evaluation results in benchmarks or downstream tasks for it are reported, it has a residual number of downloads from HF and, being based on Bloom, it inherits the restrictions from Bloom’s license and it is thus not fully open as Gervásio.
Common to these encoders other than Sabiá, which are of similar or smaller size, is that while they are publicly distributed, no public detailed presentation of them seems to be provided, be it an implementation report or a paper, either in pre-print or in peer-reviewed versions. This hampers knowing, among other aspects, which datasets were used for their training and thus hampers sensible comparison with other related work and models, which may risk being evaluated in datasets where they were trained.
@@
Looking into the collection of tasks reported to have been used to evaluate Sabiá, one finds a few that are common with the evaluation of Gervásio, namely BoolQ and WSC, from SuperGLUE, and SST2, from GLUE, which were also machine translated into PTBR to evaluate Sabiá.
Additionally, Sabiá’s authors present its performance scores in a few other downstream tasks whose datasets did not result from machine translation from English ones, but were developed originally in PTBR.
The performance scores from Sabiá’s publication are repeated in Section 5, side by side with related scores of the Gervásio PTBR, for American Portuguese. Against this background, and as it will be discussed at length in that Section, at the time of this writing and to the best of our knowledge, Gervásio offers the state of the art in terms of fully open encoders specifically improved for Portuguese in both PTPT and PTBR variants, and it is the first 7 billion parameter decoder specifically developed and distributed for the PTPT variant.
@@ 2.2.   Base models
In this connection, it is worth noting also that not only Gervásio happens to be the top performing 7 billion open encoder for Portuguese, but also that at the time of its development, it adopted the best possible setup and base code available given the goals and requirements assumed for its construction.
There is a number of multilingual encoders reported in the literature, such as mBART (Liu et al., 2020), XLM-R (Conneau et al., 2020), mT5 (Xue et al., 2021), ByT5 (Xue et al., 2022), XGLM (Lin et al., 2022), mGPT (Shliazhko et al., 2023), Bloom (Scao et al., 2022), and LLaMA (Touvron et al., 2023b), to which the promising English open models Mistral (Jiang et al., 2023) and Pythia (Biderman et al., 2023) were added in our considerations of the options available. From these possibilities, many had to be excluded given their non-open license, leaving only those from the Mistral, Bloom, Pythia, and LLaMA families as viable bases on which to build Gervásio.
@@ 3.1.   Developed datasets
To machine translate into, respectively European and American Portuguese, we resorted to DeepL, [^5] https://www.deepl.com which was set for each one of these variants, respectively.
@@
Instruction templates have been manually crafted for each task. These take the various fields in the dataset and arrange them into a prompt by, for instance, appending “Frase 1:” (Eng. “Sentence 1:”) before the first sentence of an example in the RTE dataset. These templates are listed in full detail in a dedicated webpage, [^6] https://huggingface.co/datasets/PORTULAN/extraglue of which one example is provided below in the Annex A.
@@
For testing, we reserved the translated datasets MRPC (similarity) and RTE (inference), from GLUE, and COPA (reasoning/qa), from SuperGLUE, which were taking as representatives of three major types of tasks, and were not seen during training in CLM.
@@ 4.   Models
Due to hardware limitations that imposed a shorter sequence length (512) compared to the base model (4096), instead of the typical practice of concatenating all training examples and then dividing them into batches with the same input sequence length, we separate each example individually. In other words, each example occupies the full input sequence length.
@@ Discussion
The first important result worth underling is that Gervásio largely outperforms its baseline LLaMA 2 in all tasks by both models, as reported in Table 3, except for the PTR model on the COPA task.
This demonstrates that it was rewarding to continue the causal language modeling of LLaMA 2 with the Portuguese data, even though LLaMA 2 had been pre-trained over a overwhelming majority of English data, and also despite the Portuguese dataset used to continue its pre-training being tiny (1.8 billion tokens) when compared to the one used for LLaMA 2 (2 trillion tokens).999 To further examine the outlier score of COPA in ptbr, we proceeded with cross evaluation. The PTPT model shown quite similar scores for both the PTPT and PTBR datasets, which seems to indicate that the possible cause for the outlier value did not occur with the construction of the PTBR dataset. The PTBR model, in turn, run over the PTPT testset, shown again an outlier score, similar to the outlier score obtained for PTBR, which may indicate the root of the difference occurs with the training sets. In fact, the base model LLaMA was trained on 2.8 Billion tokens of Portuguese (0.09% of the total 2 Trillion tokens used for its training in English), where PTBR texts were most probably in much superior number than PTPT ones, given the respective demographics. This indicates in which measure the two training conditions for PTBR and PTPT may differ. Nevertheless, if this larger exposure to PTBR data, by the starting model LLaMA, was the cause for the outlier value with COPA, then it will remain to expalin why the score for MRPC and RTE are in line for both PTBR and PTPT. We leave this for future research.
@@ 6.   Conclusion
This paper contributes new, instruction-tuned large language models of the encoder family of Transformers specifically developed for the Portuguese language, as well as the instruction datasets used to train and evaluate them.

### 05 Mar 2024 v2

@@ 1.   Introduction
By exploiting this approach of continuing the training of a previous strong foundation model, we contribute a new model with instruction tuning to foster the technological preparation of the Portuguese language. To the best of our knowledge, this is the first decoder under the Transformer architecture that is both (i) specifically improved for Portuguese, covering two variants of this language, namely PTBR and PTPT, and (ii) fully open, that is it cumulatively complies with all the features of being open source and openly distributed for free under a most permissive license (including for research and for commercial purposes). The model is available at https://huggingface.co/PORTULAN.
To the best of our knowledge and at the time of writing, Gervásio represents the state of the art reported in the literature for open, 7 billion parameter decoders for Portuguese, surpassing the model it is based on as well as other decoders for Portuguese of similar size. The release of Gervásio, alongside the instruction dataset used to train it and which is also a novel contribution of this paper, seeks to contribute to foster research and innovation for the language technology for Portuguese.
@@ 2.1.   Decoders for Portuguese
Looking for decoders specifically developed or improved for Portuguese that are publicly distributed and for which it is possible to find a publicly available report, to the best of our knowledge there can be found only two that, with 7 billion parameters or more, match or surpass the size of Gervásio PT* contributed in the present paper, namely the Sabiá models with 7 and 65 billion parameters (Pires et al., 2023). It is worth noting that: (i) these two models were developed for only one of the variants of Portuguese, PTBR, but not for PTPT; (ii) the 65 billion parameter model is reported in that publication but it is not distributed; and (iii) the 7 billion parameter model is distributed in a non open license, being its reuse restricted to research purposes only.
Other decoders that at the time of writing the present paper can be found of comparable size are not documented, besides being also for only one of the variants of Portuguese, namely PTBR: Boana, Cabra, Cabrita, Canarim. [^1] All on HuggingFace, at lrds-code/boana-7b-instruct, nicolasdec/CabraMistral7b-0.2, 22h/open-cabrita3b, and dominguesm/canarim-7b, respectively.
The other decoders, numbering about a dozen, that can be found for Portuguese have a smaller size, and are also only for PTBR. The largest of these, the 3 billion parameter Cabrita mentioned above, is distributed through Hugging Face (HF) and documented in a non peer-reviewed publication (Larcher et al., 2023). The second largest is Aira, [^2] On HF at nicholasKluge/Aira-2-portuguese-1B7. with 1.7 billion parameters and based on Bloom. No evaluation results on benchmarks or downstream tasks for it are reported, it has a residual number of downloads from HF and, being based on Bloom, it inherits the restrictions from Bloom’s license and it is thus not fully open as Gervásio.
Common to these decoders other than Sabiá, which are of similar or smaller size, is that while they are publicly distributed, no public detailed presentation of them seems to be provided, be it an implementation report or a paper, either in pre-print or in peer-reviewed versions. This hampers knowing, among other aspects, which datasets were used for their training and thus hampers sensible comparison with other related work and models, which may risk being evaluated in datasets where they were trained.
@@
Looking into the collection of tasks reported to have been used to evaluate Sabiá, one finds a few that are common with the evaluation of Gervásio, such as BoolQ, which were also machine translated into PTBR to evaluate Sabiá. Additionally, Sabiá’s authors present its performance scores in a few other downstream tasks whose datasets did not result from machine translation from English ones, but were developed originally in PTBR.
The performance scores from Sabiá’s publication are repeated in Section 5, side by side with related scores of the Gervásio PTBR, for American Portuguese. Against this background, and as it will be discussed at length in that Section, at the time of this writing and to the best of our knowledge, Gervásio offers the state of the art in terms of fully open decoders specifically improved for Portuguese in both PTPT and PTBR variants, and it is the first 7 billion parameter decoder specifically developed and distributed for the PTPT variant.
@@ 2.2.   Base models
In this connection, it is worth noting also that not only Gervásio happens to be the top performing 7 billion open decoder for Portuguese, but also that it adopted the best possible setup and codebase available at the time of its development given the goals and requirements assumed for its construction.
There are a number of multilingual decoders reported in the literature, such as mBART (Liu et al., 2020), XLM-R (Conneau et al., 2020), mT5 (Xue et al., 2021), ByT5 (Xue et al., 2022), XGLM (Lin et al., 2022), mGPT (Shliazhko et al., 2023), Bloom (Scao et al., 2022), and LLaMA (Touvron et al., 2023b), to which the promising English open models Mistral (Jiang et al., 2023) and Pythia (Biderman et al., 2023) were added in our considerations of the options available. From these possibilities, many had to be excluded given their non-open license, leaving only those from the Mistral, Bloom, Pythia, and LLaMA families as viable bases on which to build Gervásio.
@@ 3.1.   Developed datasets
To machine translate into European Portuguese and into American Portuguese, we resorted to DeepL, [^5] https://www.deepl.com which was set for each one of these variants, respectively.
@@
Instruction templates have been manually crafted for each task. These take the various fields in the dataset and arrange them into a prompt by, for instance, appending “Frase 1:” (Eng. “Sentence 1:”) before the first sentence of an example in the RTE dataset. A more detailed example is provided below in the Annex A.
@@
For testing, we reserved the translated datasets MRPC (similarity) and RTE (inference), from GLUE, and COPA (reasoning/qa), from SuperGLUE, which were taken as representatives of three major types of tasks, and were not seen during training in CLM.
@@ 4.   Models
Due to hardware limitations that imposed a shorter sequence length (512) compared to the base model (4096), instead of the typical practice of concatenating all training examples and then dividing them into batches with the same input sequence length, we separated each example individually. In other words, each example occupies the full input sequence length.
@@ Discussion
The first important result worth underling is that Gervásio largely outperforms its baseline LLaMA 2 in all tasks by both models, as reported in Table 3, except for the PTBR model on the COPA task.
This demonstrates that it was rewarding to continue the causal language modeling of LLaMA 2 with the Portuguese data, even though LLaMA 2 had been pre-trained over a overwhelming majority of English data, and also despite the Portuguese dataset used to continue its pre-training being tiny (1.8 billion tokens) when compared to the one used for LLaMA 2 (2 trillion tokens).888 To further examine the outlier score of COPA in ptbr, we proceeded with cross evaluation. The PTPT model shown quite similar scores for both the PTPT and PTBR datasets, which seems to indicate that the possible cause for the outlier value did not occur with the construction of the PTBR dataset. The PTBR model, in turn, run over the PTPT testset, shown again an outlier score, similar to the outlier score obtained for PTBR, which may indicate the root of the difference occurs with the training sets. In fact, the base model LLaMA was trained on 2.8 Billion tokens of Portuguese (0.09% of the total 2 Trillion tokens used for its training in English), where PTBR texts were most probably in much superior number than PTPT ones, given the respective demographics. This indicates in which measure the two training conditions for PTBR and PTPT may differ. Nevertheless, if this larger exposure to PTBR data, by the starting model LLaMA, was the cause for the outlier value with COPA, then it will remain to expalin why the score for MRPC and RTE are in line for both PTBR and PTPT. We leave this for future research.
@@ 6.   Conclusion
This paper contributes new, instruction-tuned large language models of the decoder family of Transformers specifically developed for the Portuguese language, as well as the instruction datasets used to train and evaluate them.